# Evidence for Primordial Alignment: Insights from Stellar Obliquity Measurements for Compact Sub-Saturn Systems

### v1

@@ 5 Discussion
While sub-Saturns broadly appear to boast a wide range of stellar obliquities, the few previously confirmed to be in compact multi-planet systems are exclusively aligned. Our RM measurements for two of such recently-confirmed sub-Saturns, TOI-5126 b and TOI-5398 b, continue this trend, as both are consistent with alignment. We show the updated stellar obliquity distribution for the sub-Saturn population as a function of stellar effective temperature in Figure 2. To construct this sub-Saturn sample, we combine the catalogs of Albrecht et al. (2022) and TEPCat444https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html (Southworth, 2011) accessed on March 7, 2024. We consider only RM effect measurements, which exhibit near-uniform sensitivity to the full range of possible values (compared to nonuniform methods such as starspot tracking or gravity darkening, Siegel et al. 2023; Dong & Foreman-Mackey 2023) and constitute the majority of obliquity measurements (Triaud, 2018; Albrecht et al., 2022), prioritizing those selected in Albrecht et al. (2022) (we otherwise adopt the “preferred” measurements from TEPCat). Finally, we apply a mass cut of , and select only single-star systems by cross-matching our sample with the multi-star catalog of Rice et al. (2024), which is based on the most recent data from Gaia DR3. We distinguish between sub-Saturns that are isolated and those in compact multi-planet systems, which we define as having a nearby companion with a period ratio of . All sub-Saturns considered in this work are presented in Table 2 (note that we supplement any missing parameters using the NASA Exoplanet Archive555https://exoplanetarchive.ipac.caltech.edu and values adopted by their respective reference papers). Our RM measurements allow for more reliable statistical analyses to be performed on the relationship between compact configurations and spin-orbit alignment for sub-Saturns (Section 5.1) and thus more robust constraints on the origins of spin-orbit misalignment as well as the formation histories of the sub-Saturn population (Sections 5.2 and 5.3).
@@
| TOI-5126 | b | 6253 | 1.26 | 5.46 | 21.3 | 1 | this work |
| TOI-5398 | b | 6028 | 1.105 | 10.59 | 54.4 | this work | |
@@ 5.1 Sub-Saturns in Compact Multi-planet Systems are Significantly Aligned
It is clear from Figure 2 that spin-orbit alignment is much more common for sub-Saturns in compact multi-planet systems than for isolated sub-Saturns, while no clear relation exists for the host star’s effective temperature (as seen in previous works, e.g., Albrecht et al. 2022). Our two stellar obliquity measurements increase the sample of sub-Saturns in compact configurations to five, enabling the first statistical tests of this trend. We investigate the statistical significance of preferential alignment in compact systems by comparing the number of misaligned systems found in the compact multi-planet sub-Saturn sample (zero) with that found in a series of random draws from the isolated sub-Saturn sample.
Namely, of the 22 sub-Saturns with measurements considered in this work, five are in compact multi-planet systems, while the other 17 are isolated. Of the five in compact configurations, none are misaligned, while 9/17 (53%) isolated sub-Saturns are misaligned. We define misalignment as at the level (i.e., ) and at the level (i.e., ), where we adopt the uncertainties () reported for each measurement (see Table 2). In order to perform a fair comparison between the compact multi-planet sample and the isolated sample, we randomly select and classify the alignment of five isolated sub-Saturns. We perform 100,000 iterations of these random draws, counting the number of misaligned systems for each, and fit a Gaussian function over the resulting distribution. The upper-right sub-panel of Figure 2 reveals the result, illustrating that sub-Saturns in compact multi-planet systems are more often aligned than isolated sub-Saturns to a significance level of .
@@ 5.2 Evidence for Primordial Alignment and Misalignment Via Eccentric Migration
In this work, we have demonstrated that sub-Saturns in compact planetary systems are significantly more aligned than those without nearby planetary companions (at the level). This trend of spin-orbit alignment in compact multi-planet systems around single stars is not limited to sub-Saturns, however. Short-period () Jupiters provided the first robust evidence for this phenomenon. Specifically, all measured spin-orbit angles for short-period Jupiters with nearby companions, even hot Jupiters, are consistent with alignment, e.g., Kepler-30 (Sanchis-Ojeda et al., 2012), Kepler-89 (Albrecht et al., 2013), WASP-47 (Sanchis-Ojeda et al., 2015), WASP-148 (Wang et al., 2022), TOI-1478 Rice et al. (2022), TOI-2202 (Rice et al., 2023b), TOI-1670 (Lubin et al., 2023). Conversely, isolated short-period Jupiters display a wide range of obliquities and are frequently misaligned (Albrecht et al. 2012, 2022). While there is a paucity of spin-orbit measurements for sub-Neptunes and super-Earths, most of these lower-mass planets reside in compact multi-planet systems, and those with secure obliquity measurements are aligned, e.g., Kepler-50, Kepler-65 (Chaplin et al., 2013), Kepler-25 (Albrecht et al., 2013), HD 106315 (Zhou et al., 2018; Bourrier et al., 2023), TRAPPIST-1 (Hirano et al., 2020b), TOI-1726 (Dai et al., 2020), HD 63433 (Mann et al., 2020), TOI-1136 (Dai et al., 2023), and TOI-2076 (Frazier et al., 2023).
As the spin-orbit angles of compact multi-planet systems should remain largely unaltered following the dispersal of the disk, the low stellar obliquities observed for sub-Saturns and other types of exoplanets in compact systems suggest that planetary systems are initially formed spin-orbit aligned. Furthermore, the large misalignments observed for isolated hot Jupiters and sub-Saturns are generally consistent with violent eccentric migration pathways (Dawson & Johnson, 2018; Wu et al., 2023). This scenario is additionally supported by observational evidence showing that i) both young stellar systems ( old) and planets still embedded within their debris disks are aligned (Kraus et al., 2020; Plavchan et al., 2020; Hirano et al., 2020b; Martioli et al., 2020; Palle et al., 2020; Carmichael et al., 2021; Albrecht et al., 2022; Johnson et al., 2022), and ii) that misaligned hot Jupiter systems tend to be older (Hamer & Schlaufman, 2022), reinforcing the notion that misalignments are acquired well into in the post-disk phase rather than near the onset of system formation.
@@ 5.3 On the Temperature-Obliquity Relation
There is an interesting discrepancy between the obliquity distributions of isolated hot Jupiters and isolated hot sub-Saturns: hot Jupiters are commonly misaligned around hot stars but exclusively aligned around cool stars (i.e., the relation, see Winn et al. 2010a; Albrecht et al. 2012, 2022), while hot sub-Saturns are commonly misaligned around cool stars (and no data exist for those orbiting hot stars, see Figure 2). Realignment via tidal dissipation offers a plausible explanation for this phenomenon; cool stars are susceptible to realignment by massive hot Jupiters, while lower-mass hot sub-Saturns are less likely to realign them on timescales shorter than the system’s age (e.g., see Eqn. 2 of Albrecht et al. 2012). Hence, both hot Jupiters and hot sub-Saturns may indeed become misaligned through high-eccentricity migration, but only cool stars with hot Jupiters can be realigned. This tidal realignment mechanism, however, requires finely-tuned parameters in order to completely realign all hot Jupiters around cool stars without leaving any on polar or retrograde orbits (Rogers et al., 2013; Li & Winn, 2016), which are not seen in observations.
@@ Acknowledgments
We acknowledge the independent work on the Rossiter-McLaughlin effect measurement of TOI-5398 b by Mantovan et al. Two manuscripts were prepared concurrently but independently. We appreciate the collegiality of the team regarding the manuscript submission process.
We express our gratitude for the insightful discussions with George Zhou and Chelsea Huang regarding the stellar activity of young planetary systems. We would also like to thank Armaan Goyal for insightful discussions regarding statistical analyses.
@@
M.R. acknowledges support from Heising-Simons Foundation grant 2023-4478. S.W. acknowledges support from Heising-Simons Foundation grant 2023- 4050. We acknowledge support from the NASA Exoplanets Research Program NNH23ZDA001N-XRP (grant 80NSSC24K0153). This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.

### v2

@@ 5 Discussion
While sub-Saturns broadly appear to boast a wide range of stellar obliquities, the few previously confirmed to be in compact multi-planet systems are exclusively aligned. Our RM measurements for two of such recently-confirmed sub-Saturns, TOI-5126 b and TOI-5398 b, continue this trend, as both are consistent with alignment. We show the updated stellar obliquity distribution for the sub-Saturn population as a function of stellar effective temperature in Figure 2. To construct this sub-Saturn sample, we combine the catalogs of Albrecht et al. (2022) and TEPCat666https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html (Southworth, 2011) accessed on March 7, 2024. We consider only RM effect measurements, which exhibit near-uniform sensitivity to the full range of possible values (compared to nonuniform methods such as starspot tracking or gravity darkening, Siegel et al. 2023; Dong & Foreman-Mackey 2023) and constitute the majority of obliquity measurements (Triaud, 2018; Albrecht et al., 2022), prioritizing those selected in Albrecht et al. (2022) (we otherwise adopt the “preferred” measurements from TEPCat). Finally, we apply a mass cut of , and select only single-star systems by cross-matching our sample with the multi-star catalog of Rice et al. (2024), which is based on the most recent data from Gaia DR3. This yields 22 sub-Saturn systems with measurements to comprise our statistical sample. We further distinguish between sub-Saturns that are isolated and those in compact multi-planet systems, which we define as having a nearby companion with a period ratio of . All sub-Saturns considered in this work are presented in Table 2 (note that we supplement any missing parameters using the NASA Exoplanet Archive777https://exoplanetarchive.ipac.caltech.edu and values adopted by their respective reference papers). Our RM measurements allow for more reliable statistical analyses to be performed on the relationship between compact configurations and spin-orbit alignment for sub-Saturns (Section 5.1) and thus more robust constraints on the origins of spin-orbit misalignment as well as the formation histories of the sub-Saturn population (Sections 5.2 and 5.3).
@@
| TOI-5126 | b | 6297 | 1.248 | 5.46 | 21 | 1 | this work |
| TOI-5398 | b | 6039 | 1.105 | 10.59 | 42.9 | this work | |
@@ 5.1 Sub-Saturns in Compact Multi-planet Systems are Significantly Aligned
It is clear from Figure 2 that spin-orbit alignment is much more common for sub-Saturns in compact multi-planet systems than for isolated sub-Saturns, while no clear relation exists for the host star’s effective temperature (as seen in previous works, e.g., Albrecht et al. 2022). Our two stellar obliquity measurements increase the sample of sub-Saturns in compact configurations to five, enabling the first statistical tests of this trend. We investigate the statistical significance of preferential alignment in compact systems by comparing the number of misaligned systems found in the compact multi-planet sub-Saturn sample (zero) with that found in a series of random draws from the isolated sub-Saturn sample. We perform this statistical analysis using the projected 2D spin-orbit angles , rather than the true 3D spin-orbit angles , for two reasons: i) considering solely is statistically prohibitive, as only 14 of the 22 systems in our sample have measurements, and ii) nearly all 22 systems in our sample orbit cooler stars (), which tend to have edge-on inclinations such that (see Louden et al. 2021; note that the average difference between and is just for the 14 systems with measurements for both angles).
Of the 22 sub-Saturns with measurements considered in this work, five are in compact multi-planet systems, while the other 17 are isolated. Of the five in compact configurations, none are misaligned, while 9/17 (53%) isolated sub-Saturns are misaligned. We define misalignment as at the level (i.e., ) and at the level (i.e., ), where we adopt the uncertainties () reported for each measurement (see Table 2). In order to perform a fair comparison between the compact multi-planet sample and the isolated sample, we randomly select and classify the alignment of five isolated sub-Saturns. We perform 100,000 iterations of these random draws, counting the number of misaligned systems for each, and fit a Gaussian function over the resulting distribution of systems per count. The upper-right sub-panel of Figure 2 reveals the result, illustrating that sub-Saturns in compact multi-planet systems are more often aligned than isolated sub-Saturns to a significance level of .
@@ 5.2 Evidence for Primordial Alignment and Misalignment Via Eccentric Migration
In this work, we have demonstrated that sub-Saturns in compact planetary systems are significantly more aligned than those without nearby planetary companions (at the level). This trend of spin-orbit alignment in compact multi-planet systems around single stars is not limited to sub-Saturns, however. Short-period () Jupiters provided the first robust evidence for this phenomenon. Specifically, all measured spin-orbit angles for short-period Jupiters with nearby companions, even hot Jupiters, are consistent with alignment, e.g., Kepler-30 (Sanchis-Ojeda et al., 2012), Kepler-89 (Albrecht et al., 2013), WASP-47 (Sanchis-Ojeda et al., 2015), TOI-1478 Rice et al. (2022), TOI-2202 (Rice et al., 2023b), TOI-1670 (Lubin et al., 2023). Conversely, isolated short-period Jupiters display a wide range of obliquities and are frequently misaligned (Albrecht et al. 2012, 2022). While there is a paucity of spin-orbit measurements for sub-Neptunes and super-Earths, most of these lower-mass planets reside in compact multi-planet systems, and those with secure obliquity measurements are aligned, e.g., Kepler-50, Kepler-65 (Chaplin et al., 2013), Kepler-25 (Albrecht et al., 2013), HD 106315 (Zhou et al., 2018; Bourrier et al., 2023), TRAPPIST-1 (Hirano et al., 2020b), TOI-1726 (Dai et al., 2020), HD 63433 (Mann et al., 2020), TOI-1136 (Dai et al., 2023), and TOI-2076 (Frazier et al., 2023).
As the spin-orbit angles of compact multi-planet systems should remain largely unaltered following the dispersal of the disk, the low stellar obliquities observed for sub-Saturns and other types of exoplanets in compact systems suggest that most planetary systems are initially formed spin-orbit aligned. Consequently, the dominant mechanism driving misalignment in single-star systems is likely not a universal process that operates indiscriminately across different types of systems, but instead is inherent to those with certain post-disk dynamical histories. More broadly, support for misalignment being acquired well into the post-disk phase rather than near the onset of system formation comes from evidence that i) both young stellar systems ( old) and planets still embedded within their debris disks are aligned (Kraus et al., 2020; Plavchan et al., 2020; Hirano et al., 2020a; Martioli et al., 2020; Palle et al., 2020; Albrecht et al., 2022; Johnson et al., 2022), and ii) hot Jupiter systems that are misaligned tend to be older (Hamer & Schlaufman, 2022). In combination with the large spin-orbit angles commonly observed for misaligned isolated hot Jupiters and sub-Saturns, these facts are generally consistent with violent eccentric migration as the dominant driver of misalignment (Dawson & Johnson, 2018; Wu et al., 2023); identifying the most relevant migration pathway(s), however, is beyond the scope of this work.
@@ 5.3 On the Temperature-Obliquity Relation
There is an interesting discrepancy between the obliquity distributions of isolated hot Jupiters and isolated hot sub-Saturns: hot Jupiters are commonly misaligned around hot stars () but exclusively aligned around cool stars (i.e., the relation, see Winn et al. 2010a; Albrecht et al. 2012, 2022), while hot sub-Saturns are commonly misaligned around cool stars (and no data exist for those orbiting hot stars, see Figure 2). Realignment via tidal dissipation offers a plausible explanation for this phenomenon; cool stars are susceptible to realignment by massive hot Jupiters, while lower-mass hot sub-Saturns are less likely to realign them on timescales shorter than the system’s age (e.g., see Eqn. 2 of Albrecht et al. 2012). Hence, both hot Jupiters and hot sub-Saturns may indeed become misaligned through high-eccentricity migration, but only cool stars with hot Jupiters can be realigned. This tidal realignment mechanism, however, requires finely-tuned parameters in order to completely realign all hot Jupiters around cool stars without leaving any on polar or retrograde orbits (Rogers et al., 2013; Li & Winn, 2016), which are not seen in observations.
@@ Acknowledgments
We acknowledge the independent work on the Rossiter-McLaughlin effect measurement of TOI-5398 b by Mantovan et al. We appreciate the collegiality of the team regarding the manuscript submission process.
We thank the anonymous referee for their careful review of this work and constructive feedback. We express our gratitude for the insightful discussions with George Zhou and Chelsea Huang regarding the stellar activity of young planetary systems and with Armaan Goyal regarding statistical analyses.
@@
M.R. acknowledges support from Heising-Simons Foundation grant 2023-4478. S.W. acknowledges support from Heising-Simons Foundation grant 2023- 4050. We acknowledge support from the NASA Exoplanets Research Program NNH23ZDA001N-XRP (grant 80NSSC24K0153). This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute. CP acknowledges support from ANID BASAL project FB210003, FONDECYT Regular grant 1210425, CASSACA grant CCJRF2105, and ANID+REC Convocatoria Nacional subvencion a la instalacion en la Academia convocatoria 2020 PAI77200076. The Flatiron Institute is a division of the Simons foundation.